OpenStudy (anonymous):

110 is 22% of what number? Can you show the work please?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

(x) * 0.22 = 110 ...read as 22 percent of a number = 110..all u do is rearrange. x=110/0.22 ....110 divided by 0.22 x=500 yur number is 500

OpenStudy (anonymous):

110/.22=500 so the answer is 500